OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6) y = 1 over 3x + 3 y = -3x + 3 y = 1 over 3x + 3 y = -3x + 7

OpenStudy (anonymous):

first find the slope in the equation y = -3x + 12. In y = mx + b form, the slope is in the m position...so the slope of this equation is -3. A parallel line will have the same slope. Now we use y = mx + b slope(m) = -3 (-1,6)...x = -1 and y = 6 now sub into y = mx + b and since we know the slope(m) is -3, we need to find the y intercept (b)...so we solve for b y = mx + b 6 = -3(-1) + b 6 = 3 + b 6 - 3 = b 3 = b so your parallel equation is : y = -3x + 3